# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v1

@@ Appendix E Additional Experiments
    $P_{0}:\quad(X_{1},X_{2})\sim\text{Beta}(5,5),\qquad P_{1}:\quad(X_{1},X_{2})\sim\text{Beta}(1,3).$
@@
    $\mu_{i}=\mu_{j}=:a,\qquad\sigma_{ii}=\sigma_{jj}=:b^{2},\qquad\sigma_{ij}=\sigma_{kl}=:c,$
@@
Four real-world datasets.
@@
Experiment A.1: Mismatched score function configurations.
We investigate both oracle and surrogate attack performance when using different model architectures and parameters for the score functions, with attack size $m_{a}=200$. This experiment comprises four distinct configurations for each attack type:
@@
Experiment A.2: Real-world data with NN models.
TABLE VII: Experiment A.2: FDR + NN
@@
TABLE VIII: Experiment A.2: Power + NN

### v2

@@ Appendix F Additional Experiments
$P_{0}:(X_{1},X_{2})\sim\text{Beta}(5,5),$
$P_{1}:(X_{1},X_{2})\sim\text{Beta}(1,3).$
@@
$\mu_{i}=\mu_{j}=:a,\sigma_{ii}=\sigma_{jj}=:b^{2},\sigma_{ij}=\sigma_{kl}=:c,$
@@
We consider the following four real-world datasets.
@@
Experiment A.1: Varying attack size with random forest (RF) models.
TABLE VII: Experiment A.1: FDR + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original FDR $08\pm 0.03 0.0$$8\pm 0.04 0.0$$08\pm 0.04$
oracle ($m_{a}=50$) $0.36\pm 0.02$ $0.40\pm 0.05$ $0.38\pm 0.02$
surrogate ($m_{a}=50$) $0.34\pm 0.02$ $0.20\pm 0.05$ $0.37\pm 0.02$
estimated upper bound $0.43$ $0.42$ $0.41$
oracle ($m_{a}=200$) $0.67\pm 0.00$ $0.71\pm 0.01$ $0.69\pm 0.00$
surrogate ($m_{a}=200$) $0.67\pm 0.00$ $0.64\pm 0.01$ $0.67\pm 0.00$
estimated upper bound $0.80$ $0.78$ $0.77$
TABLE VIII: Experiment A.1: Power + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original power $0.96\pm 0.02$ $0.55\pm 0.06$ $1.00\pm 0.00$
oracle ($m_{a}=50$) $0.99\pm 0.01$ $0.87\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=50$) $0.96\pm 0.02$ $0.65\pm 0.05$ $1.00\pm 0.00$
oracle ($m_{a}=200$) $0.99\pm 0.01$ $0.98\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=200$) $0.96\pm 0.02$ $0.78\pm 0.05$ $1.00\pm 0.00$
Experiment A.2: Mismatched score function configurations.
We investigate both oracle and surrogate attack performance when using different model architectures and parameters for the score functions, with attack size $m_{a}=200$. This experiment comprises distinct configurations for each attack type:
@@
TABLE X: Experiment A.3: FDR + NN
@@
TABLE XI: Experiment A.3: Power + NN
@@
Experiment A.3: Real-world data with NN models.